# From Imitation to Refinement Residual RL for Precise Visual Assembly

### v1

@@ Abstract
Behavior cloning (BC) currently stands as a dominant paradigm for learning real-world visual manipulation. However, in tasks that require locally corrective behaviors like multi-part assembly, learning robust policies purely from human demonstrations remains challenging. Reinforcement learning (RL) can mitigate these limitations by allowing policies to acquire locally corrective behaviors through task reward supervision and exploration. This paper explores the use of RL fine-tuning to improve upon BC-trained policies in precise manipulation tasks. We analyze and overcome technical challenges associated with using RL to directly train policy networks that incorporate modern architectural components like diffusion models and action chunking. We propose training residual policies on top of frozen BC-trained diffusion models using standard policy gradient methods and sparse rewards, an approach we call ResiP (Residual for Precise manipulation). Our experimental results demonstrate that this residual learning framework can significantly improve success rates beyond the base BC-trained models in high-precision assembly tasks by learning corrective actions. We also show that by combining ResiP with teacher-student distillation and visual domain randomization, our method can enable learning real-world policies for robotic assembly directly from RGB images. Find videos and code at https://residual-assembly.github.io.
Keywords: Residual Learning, Robotic Assembly, Combining BC and RL
1 Introduction
Despite significant advancements in robot learning, autonomous systems still struggle with manipulation tasks requiring high precision, robustness to a wide array of initial conditions, and operation directly from RGB sensors. An example of such a challenging scenario is autonomous assembly, where a robot is presented with a set of multiple parts to pick up and precisely fit together. Such systems typically utilize engineered fixtures for consistent initial part positioning, making assembly robots expensive and impractical to adapt across diverse tasks. Learning to perform these tasks from vision can improve the adaptability and robustness of such systems, but the high-precision, contact-rich nature of such tasks poses substantial challenges for current robotic learning frameworks.
Behavior cloning (BC), has re-emerged as a popular technique for acquiring manipulation control policies [1, 2, 3, 4, 5, 6]. This is largely due to the ease of directly specifying complex tasks by providing demonstrations. However, BC requires large amounts of costly demonstration data to enable generalization to different scene configurations [5]. BC-trained policies are also often limited by the nature of human demonstrations that directly achieve the task. This can mean such policies never learn to perform corrective actions that recover from deviations that inevitably occur during deployment. Reinforcement learning (RL) offers a complementary paradigm, wherein locally corrective behavior can be learned by exploring and discovering action sequences that lead to downstream task success.
Although RL offers appealing advantages, many of its successes in robotics depend on carefully crafted reward functions that may sacrifice generality across tasks [7, 8, 9]. A common mitigation is to bootstrap the RL process by starting with a small set of demonstrations [10, 11, 12] – by initializing with a policy that occasionally solves the task, supervision from simple sparse rewards can be enough for RL to progress effectively. However, naively trained BC policies may fail to achieve any meaningful success with a small number of demonstrations in complex tasks like assembly [13]. Innovations in policy design, such as diffusion models [14] and action chunking [6], can alleviate some of this difficulty and enable non-zero success rates for challenging tasks using a modest number of demonstrations [13, 15]. However, introducing these modern architectural components makes it difficult to perform fine-tuning with RL. For instance, training diffusion with RL can require complex reformulation of the de-noising process to enable policy gradient estimation [16]. We also find that action chunking tends to make direct RL fine-tuning harder and unstable due to the expanded action space dimension induced by chunking.
In this paper, we propose adaptations to the BC + RL pipeline that facilitate its application to complex tasks like furniture assembly [15]. Our key contribution is to enable RL fine-tuning to work with modern advancements like diffusion policies and chunked action predictions. We first provide an analysis and empirical evidence supporting the need for action chunking and expressive policy architectures for achieving success rates that are high enough to bootstrap RL. We then propose training residual policies [17, 18, 19] with on-policy RL and sparse task rewards to improve upon pure BC, an approach we term ResiP (Residual for Precise manipulation). These residual models learn to predict corrective actions that more reliably achieve high precision. In this way, we side-step the aforementioned challenges of directly altering the base policy with RL, while improving overall task success. Finally, by distilling behaviors learned with RL into a large synthetic dataset of trajectories and combining them with a limited set of real-world demonstrations, we showcase the benefits of our residual learning approach, ResiP, in enabling real-world assembly policies that operate directly from RGB images.
2 Problem Setup and Approach Overview
Our goal is to develop a pipeline that enables robots to perform precise manipulation tasks without requiring massive amounts of manual human effort. While our pipeline can be applied to many tasks, we focus on applications to the domain of multi-step assembly directly from RGB images. Here, we introduce the specific assumptions of our problem setup and detail the main ideas of the pipeline that we propose.
@@ 2 Problem Setup and Approach Overview
We assume an assembly task is specified to the robot via a mid-sized ($\sim$50-90) set of expert demonstrations. The tasks each have an underlying success criterion depending on a set of required alignments between the parts. Each task consists of a fixed set of rigid object parts for which we have access to CAD models, such that both the parts and the entire assembly process can be accurately simulated. Each task may have long horizons ($\sim$750-1000 steps) and require sequencing of behaviors such as corner alignment, 6-DoF grasping, reorientation, insertion, and screwing. Multi-part assembly interactions are simulated using the SDF-based collision geometry representations featured in the Factory [20] extension of NVIDIA’s Isaac Gym simulator. We use the tasks one_leg, round_table, and lamp from the FurnitureBench [15] task suite.
@@
Preliminaries. We formulate the robot’s task as a discrete-time sequential decision-making problem. In each time step $t$, the robot receives either an observation $o_{t}\in\mathcal{O}$ if it operates in the real world or the state of the system $s_{t}\in\mathcal{S}$ if it operates in simulation. After receiving it, the agent produces an action $a_{t}$ to execute in the environment. The action space in both simulation and the real world is the desired end-effector pose $\mathbf{T}{}^{\text{des}}\in$ $\text{SE}(3)$. We use a differential inverse kinematics controller to convert the desired end-effector pose commands into joint position targets, which are tracked with a low-level PD controller. The real world observation space $\mathcal{O}$ contains the robot end-effector pose $\mathbf{T}{}\in$ $\text{SE}(3)$, robot end-effector spatial velocity $\mathbf{V}\in\mathbb{R}^{6}$, the gripper width $w_{\text{g}}$, and RGB images from a fixed front-view camera ($I^{\text{front}}\in\mathbb{R}^{h\times w\times 3}$) and a wrist-mounted camera ($I^{\text{wrist}}\in\mathbb{R}^{h\times w\times 3}$), each with unknown camera poses. In the simulated task variants, the system states space $\mathcal{S}$ contains the same end-effector pose $\mathbf{T}$, spatial velocity $\mathbf{V}$, and gripper width $w_{\text{g}}$, along with the 6-DoF poses of all the parts in the environment $\{\mathbf{T}{}^{\text{part}_{i}}\}_{i=1}^{\text{num\_parts}}$. In addition to the state/observation, the simulated agent receives a binary reward signal from the environment, indicating whether two parts of the assembly have achieved their required geometric alignment.
BC + RL Teacher-Student Distillation Pipeline
The pipeline’s high-level goal is to leverage a small amount of human demonstrations and the availability of a simulator to develop a vision-based manipulation policy. To this end, the pipeline consists of two phases:
  • Phase one: We train an agent to perform the task in a simulated environment using RL. To overcome the need for manual reward engineering, we start with an initial policy pretrained with behavior cloning on a limited set of demonstrations collected in the simulation. We then fine-tune the policy with RL, with only a sparse reward signal provided upon successful task completion.
  • Phase two: Next, we distill the fine-tuned RL expert policy into a dataset of high-quality RGB image observations and corresponding actions for vision-based policy distillation (using BC). This dataset is combined with a small set of real-world task demonstrations, which helps bridge the simulation-to-reality gap for the final vision-based manipulation policy.
3 Methods
Figure 1: Pipeline overview. (1) Beginning with a policy trained with BC in simulation, (2) we train residual policies to improve task success rates with RL and sparse rewards. (3) We then distill the resulting behaviors to a policy that operates on RGB images. (4) By combining synthetic data with a small set of real demonstrations, (5) we deploy assembly policies that operate from RGB images in the real world.
3.1 Imitation Learning
The first step in the pipeline is to use a small number of human demonstrations to train a “base” policy that serves as the starting point for RL training. For each task, we collect a dataset of 50 demonstrations in simulation, $\mathcal{D}_{\text{sim}}$. This dataset contains trajectories, e.g., $\mathcal{D}=\{\tau_{1},...,\tau_{N}\}$, where each trajectory contains the system states $s_{t}$, and robot actions $a_{t}$, i.e., $\tau_{i}=\{(s_{t},a_{1}),...,(s_{T},a_{T})\}$, with $T$ being the trajectory length. We obtain these trajectories by teleoperating the simulated robot to complete the task.
Using the simulation dataset $\mathcal{D}_{\text{sim}}$, we train the base policy $\pi_{\text{base}}$ with Behavior Cloning (BC), i.e., we maximize the likelihood of the data by optimizing $\max_{\pi}\ \mathbb{E}_{(a_{t},s_{t})\sim\mathcal{D}_{\text{sim}}}\left[\log\pi(a_{t}|s_{t})\right]$. We use Diffusion Policy [14] as the base model, which has shown strong empirical performance in handling difficult manipulation tasks with relatively small datasets [13]. Consistent with recent advancements [14, 6, 13], our policy framework enhances its performance by predicting multiple future actions in chunks, as opposed to individual actions at each timestep. We denote the length of future action sequences predicted by the policy as $T_{a}$, the output as $\mathbf{a_{t}}=[a_{t},...,a_{t+T_{a}}]$. When predicting an action chunk $\mathbf{a_{t}}$ of length $T_{a}$, we only execute a subset $[a_{t},...,a_{t+T_{\text{exec}}}]$, with execution horizon $T_{\text{exec}}\leq T_{a}$.
3.2 Online Reinforcement Learning with Residual Policies
Given the initial base policy $\pi_{\text{base}}$, we aim to improve it using RL. However, directly fine-tuning diffusion models with RL is an active area of research [21, 22, 16, 23], made difficult because of the multi-step inference process and unavailability of the policy action log-probabilities. Another category of methods upweight high-quality model outputs via importance sampling, return conditioning, or augmenting the original de-noising objective with a loss term for maximizing a Q-function [24, 25, 26, 27, 28]. However, these methods mainly enable better extraction of high-quality behavior in the data, while we are more concerned with learning new, corrective behaviors. Action chunks can also make optimization with policy gradients more difficult. This is partly due to chunking increasing the effective action space (e.g., chunks of 8 actions increase the dimension 8$\times$), which we find to increase RL training instability (see Sec. 4.1). This means fine-tuning other popular BC architectures, like the Action-Chunked Transformer (ACT) [6], also brings technical challenges. Furthermore, recent work has shown that fine-tuning large pre-trained models can lead to forgetting of capabilities if the agent does not visit states seen during pre-training frequently enough [29].
Figure 2: Per-timestep residual policies trained with PPO to locally correct action chunks predicted by a BC-trained diffusion policy.
We side-step these complications by training a residual Gaussian MLP policy $\pi_{\text{res}}$. This policy takes as input both the system state and the action predicted by the diffusion policy $\pi_{\text{base}}$ and produces an “action correction” that modifies the action as $a_{t}=a_{t}^{\text{base}}+\alpha\cdot a_{t}^{\text{res}}$, with $\alpha\leq 1$ being a coefficient controlling what scale the residual policy operates on. We denote the resulting combined policy $\boldsymbol{\pi}$. This decoupling has several advantages. First, we can regard the diffusion model as a black box and not change its parameters [17] [^1] This decoupling also implies that the ACT or any other strong BC or offline RL method can straightforwardly serve as the base model without changing anything else in the pipeline, an advantage of residual policy learning emphasized by other prior work in robotics [17, 18].. This also allows the use of different prediction horizons $T_{a}$ for the two policies. This flexibility in action horizon is helpful as most RL algorithms optimize single-action policies. It also removes the need to explicitly regularize the fine-tuned policy to stay near the pre-trained policy, which is often necessary to achieve stable optimization [30, 31].
Given the above, we train the $\pi_{\text{res}}$ using standard PPO [32] with the alternation that we augment the state space with the base action. The base model observes the current state $s_{t}$ and outputs a chunk of actions $\mathbf{a_{t}}{}$. For each action in the chunk, we concatenate it with the current observation $s_{t+i}^{\text{res}}=[s_{t+i},a^{\text{base}}_{t+i}]$, and predict the correction $a_{t+i}^{\text{res}}\sim\pi^{\text{res}}(\cdot|s_{t+i}^{\text{res}})$ for $i=0,...,T_{a}-1$. Besides the technical convenience, one can look at residual policy as a way to incorporate an inductive bias that the policy should only learn locally corrective actions [18]. This is motivated by our qualitative observations that most of the base BC policy failures are due to slight imprecision near “bottleneck” transitions when performing skills like grasping and insertion (see Fig. 3).
3.3 Distillation for RGB-based Visual Policy
Once the RL training has converged, the policy typically exhibits significantly higher success rates compared to the initial base policy. Our objective is to distill this enhanced performance from the state-based policy into a vision-based policy that operates solely on RGB images. Following the established teacher-student distillation paradigm [9, 33, 7], we generate a dataset of successful trajectories, $\mathcal{D}_{\text{synth}}$, where the environment’s observations, $o_{t}$, replace the states, i.e., $\mathcal{D}_{\text{synth}}=\{\tau_{\text{synth},1},...,\tau_{\text{synth},N}\}$ and $\tau_{\text{synth},i}=\{(o_{1},a_{1}),...,(o_{T},a_{T})\}$. We utilize this dataset to train a final image-conditioned diffusion policy with BC.
For real-world transfer, we enhance the synthetic dataset $\mathcal{D}_{\text{synth}}$ by re-rendering its trajectories in Isaac Sim. This process improves image quality and introduces variability in environmental conditions such as object and table colors, textures, lighting, and camera perspectives, many of which cannot easily be done with standard image augmentation techniques. We denote this refined dataset as $\mathcal{D}_{\text{synth-render}}$. To ease the difficulty of zero-shot sim-to-real with RGB images, we opt for a co-training approach, integrating the synthetic dataset with a small set of real-world task demonstrations, $\mathcal{D}_{\text{real}}$, which similarly comprises only environmental observations without ground truth poses. This combined dataset, $\mathcal{D}_{\text{real}}\cup\mathcal{D}_{\text{synth-render}}$, is used to train the final student policy through BC. While loss reweighting or upsampling may be helpful given the imbalanced dataset sizes, we find that simply mixing the two datasets still provides performance improvement beyond training purely with real data (see Fig. 2).
4 Experiments and Results
In Sec. 4.1, we investigate the requirement of complex BC methods and the impact of our residual reinforcement learning approach on improving the success rates of policies trained with imitation learning. Next, in Sec. 4.2, we study the performance of downstream distillation from synthetic RL data, examining the relationship between the quantity and quality of the data and the resulting performance of the distilled vision-based policies. Finally, in Sec. 4.3 we demonstrate our approach enabling precise assembly tasks on a physical robot directly from vision.
4.1 Improving Imitation Learning with Online Residual Learning
Figure 3: Representative failures from BC policies. RL-trained residual policies dramatically improve succcess rate by correcting for these errors.
Why Action Chunking and Diffusion Policies?
Simple feed-forward MLPs of modest size have shown impressive performance in many domains when trained with RL [7, 33, 34], and offer a natural starting point for RL fine-tuning after BC pre-training. However, we find that MLP policies trained to predict single actions (MLP-S) completely fail across all tasks we consider. Therefore, we also trained MLP policies with action chunking (MLP-C). When we introduce chunking, MLP performance improves drastically. However, we also find that the more complex diffusion policy architecture generally outperforms MLPs, especially in tasks of intermediate difficulty. For example, an improvement from 10% success rate to 29% for the one_leg task on medium randomness, making subsequent fine-tuning far easier. In one case, lamp on low randomness, MLP-C outperformed DP. In qualitative evaluations, we find that DP has smoother and faster actions, which generally is beneficial, but in this case, seems to hurt performance, as it tends to retract before the gripper closes. We also find that all methods struggle for the hardest tasks, i.e., on which MLP-C and DP both achieve less than 5% success rate, indicating that there is still a need for stronger BC methods.
Online Policy Fine-Tuning Comparison
Next, we investigate the effectiveness of our residual RL approach, ResiP, in improving the performance of policies trained with BC. We train residual policies using PPO [32] on top of the diffusion policy and compare it to two baselines – (1) directly fine-tune the MLP-C with PPO while regarding every action chunk as one concatenated action. (2) Using the recently proposed IDQL algorithm [24] where one samples multiple times from the diffusion policy and chooses which action chunk to execute based on its Q-value. We have found that learning a good Q-value estimator from offline data alone, as proposed in IDQL, is challenging with only 50 demonstrations. Therefore, we learn it in on-policy matter similar to [35].
The results in Tab. 1 show that our residual RL approach significantly improves success rates over BC and outperforms alternative RL fine-tuning methods. For tasks with lower initial randomization, such as the one_leg task, the residual policy increases the success rate from 55% to 95%. However, for tasks with higher part randomization, we observe performance saturates at lower success rates. We hypothesize this is because the residual policy is limited by the performance of the base model.
@@ 4.1 Improving Imitation Learning with Online Residual Learning
Methods One Leg Round Table Lamp
Low Med Low Med Low Med
IL
MLP-S 0 0 0 0 0 0
MLP-C 45 10 5 2 8 1
DP 54 29 12 4 7 2
RL
MLP-C + PPO 70 28 38 6 32 2
DP + IDQL 52 13 18 3 11 1
ResiP (ours) 95 68 55 10 60 8

### v3

# From Imitation to Refinement – Residual RL for Precise Assembly

@@ Abstract
Advances in behavior cloning (BC), like action-chunking and diffusion, have enabled impressive capabilities. Still, imitation alone remains insufficient for learning reliable policies for tasks requiring precise aligning and inserting of objects, like assembly. Our key insight is that chunked BC policies effectively function as trajectory planners, enabling long-horizon tasks. Conversely, as they execute action chunks open-loop, they lack the fine-grained reactivity necessary for reliable execution. Further, we find that the performance of BC policies saturates despite increasing data. Reinforcement learning (RL) is a natural way to overcome BC’s limitations, but it is not straightforward to apply directly to action-chunked models like diffusion policies. We present a simple yet effective method, ResiP (Residual for Precise Manipulation), that sidesteps these challenges by augmenting a frozen, chunked BC model with a fully closed-loop residual policy trained with RL. The residual policy is trained via on-policy RL, addressing distribution shifts and introducing reactive control without altering the BC trajectory planner. Evaluation on high-precision manipulation tasks demonstrates strong performance of ResiP over BC methods and direct RL fine-tuning.
Videos, code, and data are available at https://residual-assembly.github.io.
I Introduction
Robotic manipulation tasks that require both long-horizon planning and high-precision control, such as assembly, remain a significant challenge in robotics [2, 3, 4, 5, 6]. As an example, consider the furniture assembly task depicted in LABEL:fig:overview (Left), where a robot performs assembly by sequentially grasping and then re-orienting a table leg into the correct pose, performing a precise insertion, and finally screwing the leg in place. This representative task involves executing a specific sequence of skills over hundreds of timesteps, with each stage dependent on the successful completion of the previous. Small imprecisions anywhere in the chain can lead to task failure, underscoring the need for reliable execution of diverse skills in the sequence.
Behavior cloning (BC) is a popular approach for teaching robots various manipulation skills [7, 8, 9, 10, 11, 12, 13, 14, 6, 15]. Recent innovations in BC, such as diffusion models [16, 17, 18, 19] and action chunking [20, 6, 16, 18], have improved learning long-horizon, complex behaviors from demonstrations [21, 15]. However, our analysis shows fundamental limitations in BC when applied to tasks requiring high precision. For example, in the insertion task shown in LABEL:fig:overview (Left), a diffusion-based BC model achieves a success rate of $\sim$50with a handful of demonstrations, but performance plateaus even when scaling up to large datasets (e.g., 100,000 demonstrations), reaching only about $\sim$80success, as seen in LABEL:fig:overview (Right). Other recent work finds similar performance saturation [22, 23, 24].
We hypothesize that this saturation occurs because of two issues: First, BC methods suffer from compounding errors due to distributional shifts; the policy operates on states that deviate from those seen during training, leading to errors that accumulate over time [22]. While approaches like Dataset Aggregation (DAgger) [1] can mitigate this error mode by collecting on-policy data, such approaches assume an expert who can be queried on demand, which is usually unavailable. Second, modern BC policies, particularly ones employing action chunking, act more like open-loop “planners” than reactive controllers. Chunked BC policies are limited in compensating for noise and disturbances (i.e., non-reactive) because each chunk runs in an open loop. In tasks such as furniture assembly, specific “bottleneck” states, like insertions, require precise actions at specific time steps. If these critical moments fall within an action chunk, the BC trajectory planner cannot make real-time adjustments to compensate for inaccuracies in the planned actions. This inability to correct actions can lead to task failure in high-precision tasks.
Reinforcement Learning (RL) promises to autonomously learn corrective behaviors through interaction and exploration guided by a reward function [25, 26]. RL has been effective in training robust reactive controllers that can handle a wide array of tasks [27, 28, 29, 30, 31]. However, applying RL to solve complex, long-horizon, and precise tasks like robotic assembly requires extensive reward engineering that can be prohibitively difficult and time-consuming [32]. RL requires vast data, which is costly to collect, particularly in real-world settings.
We present a way to leverage the strengths of BC and RL that reduces human effort in obtaining reliable policies by building on recent advances in BC methods and sim-to-real RL training. BC learns a “planner” that predicts trajectory segments (i.e., an action chunk executed open loop) conditioned on the agent’s observations from a small set of demonstrations, providing a skeleton for RL training. RL augments the BC-learned planner with a closed-loop controller capable of fine-grained corrections that mitigate the limitations of open-loop action chunk execution and collects on-policy data to ensure reliable plan execution. While combining BC and RL has a long history [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45], the recent advancements in BC architectures present new challenges for fine-tuning the learned policy with RL. The structure of diffusion models (iterative refinement) and action-chunked policies (large effective action spaces) complicate the application of standard RL algorithms, often leading to instability or requiring significant architectural modifications [46, 15, 47].
We propose a simple yet effective method to combine the strengths of modern chunked BC and on-policy RL controllers, called ResiP (Residual for Precise Manipulation), shown in Fig. 1. Our key insight is recognizing that modern BC policies function as trajectory planners that generate action sequences but crucially lack the fine-grained reactivity and corrections required for reliable execution. We introduce closed-loop corrections by augmenting a frozen, chunked BC model with a small, single-step residual policy [48] trained via on-policy RL. The residual policy observes each state and action and predicts a corrective action to be added to the action predicted by the BC model. This design sidesteps the difficulties of fine-tuning complex BC models with RL while addressing BC’s limitations. First, the residual RL policy operates on-policy, learning reactive behaviors directly through interaction with the environment, thus overcoming issues of distributional shift. Second, providing closed-loop control enables real-time adjustments during task execution, improving robustness and precision, especially in bottleneck stages like insertions. Finally, the residual training is supervised with a sparse task-completion reward, avoiding reward engineering.
We evaluate ResiP on a range of high-precision manipulation tasks from the challenging FurnitureBench [5] and Factory [49] benchmarks and two hand-designed tasks. Our experiments demonstrate that ResiP outperforms pure BC methods and traditional RL fine-tuning, for example, achieving success rates up to $\sim$98starting from 50 demonstrations for the one_leg task. At the same time, increasing data for BC saturates at $\sim$80success with 100K demonstrations for the same task. We also find that ResiP, combined with straightforward applications of teacher-student distillation, visual domain randomization, and co-training, we can transfer policies from simulation to physical robot hardware using RGB observations.
Figure 1: Overview of ResiP. A pretrained chunked base policy predicts an action chunk of $T_{a}$ future actions. For every timestep, the residual model observes the current state $s_{t}$ and predicted base action $a_{t}^{\text{base}}$ and corrects.
II Problem Setup and Approach Overview
Figure 2: Initial and goal states for our 6 tasks. The first three tasks, one_leg, round_table, and lamp, come from the FurnitureBench [5] task suite. mug-rack we created by scanning objects and importing them into the simulation. peg-in-hole is from the Factory [49] task suite. biman-insert we created using the simulation-based teleoperation system DART [50]. Our task suite exhibits diverse challenges like long horizons, tight tolerances, bimanual control, and multi-modal success criteria.
@@ II Problem Setup and Approach Overview
Our method applies to tasks that can be simulated using already available assets (e.g., CAD models) or using assets that can be obtained via scans of real-world environments [30, 51]. While our method is general, we empirically verify it on a set of challenging assembly tasks of one_leg, round_table, and lamp taken from the FurnitureBench [5] task suite, the peg-in-hole task from [49], a custom mug-hanging task we call mug-rack, and a custom bimanual precise insertion task we call biman-insert. Visualizations of initial state distributions and detailed task descriptions are provided in App. B, and task rollouts can be seen on the project website. Multi-part assembly interactions are simulated using the SDF-based collision geometry representations provided as part of the Factory [49] extension of NVIDIA’s Isaac Gym simulator [52]. The biman-insert task is implemented using the MuJoCo simulator [53], and demos are provided using the DART teleoperation system from [50].
We define a sparse task-completion reward set to 1 for each task when a pair of parts has been fully assembled and 0 otherwise. Many of our tasks are long horizon (up to $\sim$750-1000 steps at 10Hz) and require sequencing of behaviors such as 6-Degree-of-Freedom (DoF) grasping, reorientation, insertion, and screwing (see, e.g., LABEL:fig:overview (Left)).
Preliminaries
Each task is formulated as a discrete-time sequential decision-making problem. At every time-step $t$, the robot receives an observation $o_{t}\in\mathcal{O}$ corresponding to the world and robot state $s_{t}\in\mathcal{S}$ (available only in simulation) and uses it to predict an action $a_{t}$ that is then executed.
The policy operates at 10Hz on a 7-DoF Franka Emika Panda robot arm for all tasks but biman-insert, which operates at 50Hz on two Franka Panda arms. The action space consists of the desired end-effector pose $\mathbf{T}{}^{\text{des}}\in$ $\text{SE}(3)$, which includes both position and orientation and a binary gripper command for opening/closing the parallel-jaw gripper. These desired end-effector poses are converted to joint position targets using differential inverse kinematics [54], then tracked using a low-level PD controller running at 1KHz with manually specified stiffness and dampening parameters.
@@
In the simulated task variants, the system state space $\mathcal{S}$ contains the same end-effector pose $\mathbf{T}$, spatial velocity $\mathbf{V}$, and gripper width $w_{\text{g}}$, along with the 6-DoF poses of all the parts in the environment $\{\mathbf{T}{}^{\text{part}_{i}}\}_{i=1}^{\text{num\_parts}}$. In addition to the state and observation, the simulated agent receives a binary reward signal, indicating whether a pair of assembly parts has achieved their required geometric alignment. For example, in the lamp task, the policy receives two binary rewards: the first when the bulb is fully screwed in and the second when the lamp shade is successfully placed. The real world observation space $\mathcal{O}$ contains the robot end-effector pose $\mathbf{T}{}\in$ $\text{SE}(3)$, robot end-effector spatial velocity $\mathbf{V}\in\mathbb{R}^{6}$, the gripper width $w_{\text{g}}$, and RGB images from a fixed front-view camera ($I^{\text{front}}\in\mathbb{R}^{h\times w\times 3}$) and a wrist-mounted camera ($I^{\text{wrist}}\in\mathbb{R}^{h\times w\times 3}$), each with unknown camera poses.
@@
Methods one_leg round_table lamp mug-rack peg-in-hole biman-insert
Low Med Low Med Low Med Low Low Low
IL MLP-S 0 0 0 0 0 0 0 2 0
MLP-C 45 10 5 2 8 1 21 2 7
DP 54 29 12 4 7 2 26 5 33
RL MLP-C + PPO 70 28 38 6 32 2 23 4 30
DP + IDQL 52 27 18 3 11 1 31 3 40
ResiP (ours) 98 76 94 77 97 70 88 99 93